OpenStudy (anonymous):

A calculator will be helpful in solving the following problem. After receiving an inheritance of $11,600, a man deposits the money in an account paying an annual rate of 7.2%, compounded daily. How much money will be in the account at the end of 4 years? (Round your answer to the nearest cent.)

jimthompson5910 (jim_thompson5910):

A=P(1+r/n)^(n*t) A=11600(1+0.072/365)^(365*4) A=11600(1+0.000197260273972603)^(365*4) A=11600(1.00019726027397)^(365*4) A=11600(1.00019726027397)^(1460) A=11600(1.33371942362715) A=15471.145314075 A=15471.15 So $15,471.15 will be in the account at the end of 4 years

OpenStudy (anonymous):

A= 11600(1+ 0.072/365)^(365*4) = 15471.15